OpenStudy (anonymous):

the carpenters built 28 tables 19 shelves and 33 cabinets one work week if they work 8 hrs per day how many items did they build on an average day? explain

hero (hero):

Divide each thing by eight and round down then add it all up

hero (hero):

Or you can add everything up then divide by 8

hero (hero):

So (28 + 19 + 33)/8 = 80/8 = 10 So in eight days the built a total of 80 items. If we divide 80 by 8, we get 10 items which equals the amount they can build in one day.